OpenStudy (dillonmerv):

Suppose f(x) is such that f''(x) is continuous on [0, π], f(π) = 1 and {Integral with upper bound π and lower bound 0 (f(x) + f''(x)) sin(x)dx = 2.} Find f(0). I know that I should break the integral into two separate integrals as in: {(Integral with upper bound π and lower bound 0) f(x) sin(x)dx + (Integral with upper bound π and lower bound 0) f''(x) sin(x)dx. Don't know what to do next.

OpenStudy (tylerd):

havent done math in awhile, ill take a crack at it though

OpenStudy (tylerd):

what section of diffy q is this

OpenStudy (dillonmerv):

calc 2

OpenStudy (zzr0ck3r):

\(\int uv'dx=uv-\int u'vdx\) \(\int_0^\pi f(x)\sin(x)dx+\int_0^\pi f^{''}(x)\sin(x)dx=2\) On the first integral we let \(v=-\cos(x)\) and \(u=f(x)\) \(\int_0^\pi f(x)\sin(x)dx=\\ \int_0^\pi f(x)(\cos(x))'dx=\\-f(x)\cos(x)|_o^\pi-\int_0^\pi -f(x)\cos(x)dx=\\-1(-1)+f(0)+\int_0^\pi f^{'}(x)\cos(x)dx \) Do something similar on the second integral (let \(v=f^{'}(x)\) and \(u=\sin(x)\)) and get \(\int_0^\pi \sin(x)f^{''}(x)dx=-\int_0^\pi \cos(x)f^{'}(x)dx\) Conclude \(f(0) = 1\)

OpenStudy (dillonmerv):

Looks very good. Thanks!

OpenStudy (dillonmerv):

If I have any issues I'll get back to you.

OpenStudy (zzr0ck3r):

Sounds good. I might be sleeping, but someone should be around.